# Point-hyperplane incidences via extremal graph theory

### v1

@@ 4 Large complete bipartite subgraphs in incidence graph
The last inequality is true due to simple calculations. We should also verify that $G^{\prime}$ satisfies the conditions of Lemma 4.3 by showing that $\varepsilon^{\prime d-1}n^{\prime}\geq 100^{d-1}$ and $\varepsilon^{\prime d}m^{\prime}\geq 100^{d}$. But this also follows, since we have already shown that $\varepsilon^{\prime d-1}n^{\prime}m^{\prime}\geq\varepsilon^{d-1}mn$, which implies both $\varepsilon^{\prime d-1}n^{\prime}\geq\varepsilon^{d-1}n\geq 100^{d-1}$ and $\varepsilon^{\prime d}m^{\prime}\geq\varepsilon^{d}m\geq 100^{d}$ (since $\varepsilon^{\prime d-1}n^{\prime}m^{\prime}\geq\varepsilon^{d-1}mn$ also implies $\varepsilon^{\prime}\geq\varepsilon$). We conclude that $\operatorname{rs}(G)\geq(\varepsilon/100)^{d-1}mn$, as promised.
In what follows, we assume that $G$ contains no vertex $v\in A$ of degree less than $\left(1-\frac{2}{d}\right)\varepsilon m$ and no vertex $v\in B$ of degree less than $\left(1-\frac{2}{d}\right)\varepsilon n$. Let us now focus on the main argument. Recall that we proceed by induction on $d$, where the case $d=3$ was discussed in Section 2 as Lemma 2.2.
@@
We choose a good $(d-1)$-tuple $(v_{1},\dots,v_{d-1})$ with the maximal size of the common neighbourhood, which we denote by $X=N(v_{1},\dots,v_{d-1})$. By Lemma 4.5, there must be a good $(d-1)$-tuple with $|X|\geq\left(\frac{\varepsilon}{6}\right)^{d-1}|B|$. If $A$ does not contain a good $d$-tuple in $G$, then any vertex $v\in A$ either has $X\subseteq N(v)$ or $|N(v)\cap X|=1$. Lemma 4.4 applied with $\gamma=(\frac{\varepsilon}{6})^{d-2}$ shows that $|X|\leq\left(\frac{\varepsilon}{6}\right)^{d-2}|B|$. We conclude that one can partition $A$ into two sets, $V_{1}$ and $V_{2}$, where $V_{1}=\{v\in A:X\subseteq N(v)\}$ and $V_{2}=\{v\in A:|X\cap N(v)|\leq 1\}$.
@@
Let us consider the edges incident to $X$. By our assumption on the minimum degree of vertices of $X$, we have $e(X,A)\geq|X|\cdot\left(1-\frac{2}{d}\right)\varepsilon n\geq\frac{\varepsilon}{3}n|X|$. On the other hand,
$\frac{\varepsilon}{3}n|X|\leq e(X,A)=e(X,V_{1})+e(X,V_{2})\leq|X||V_{1}|+|V_{2}|\leq|X||V_{1}|+n.$
@@
Since $|X|\geq\left(\frac{\varepsilon}{6}\right)^{d-1}m$ and $(\frac{\varepsilon}{100})^{d}m\geq 1$, we have $|X|\geq\frac{12}{\varepsilon}$. Thus, $\frac{\varepsilon}{12}n|X|\geq n$ and so $\frac{\varepsilon}{4}n|X|\leq|X||V_{1}|$, showing that $|V_{1}|\geq\frac{\varepsilon}{4}n$.
@@
since $6^{d-2}\geq 2d$ for all $d\geq 3$. Thus, the sizes of the vertex classes of $G^{\prime}$ are $n^{\prime}=|V_{1}|\geq\frac{\varepsilon n}{4}$ and $m^{\prime}=|B\backslash X|\geq(1-\frac{5/2}{d})m$. Finally, the above computation shows that $e(G^{\prime})=\varepsilon^{\prime}m^{\prime}n^{\prime}\geq(1-\frac{5/2}{d})\varepsilon m^{\prime}n^{\prime}$.
Next, our plan is to apply the induction hypothesis for $d-1$ on the graph $G^{\prime}$ to find a large complete bipartite subgraph. Let us first argue that $G^{\prime}$ does not have a good $(d-1)$-tuple. Suppose for contradiction that $G^{\prime}$ contains a good $(d-1)$-tuple $u_{1},\dots,u_{d-1}$. The vertices $u_{1},\dots,u_{d-1}$ form a good $(d-1)$-tuple in the graph $G$ as well, and their common neighbourhood contains both $X$ and at least two vertices from $B\backslash X$ (since this is a good $(d-1)$-tuple in $G^{\prime}$). But this implies $u_{1},\dots,u_{d-1}$ is a $(d-1)$-tuple with a larger common neighbourhood in $G$ than $v_{1},\dots v_{d-1}$, contradicting the maximality. Thus, we conclude that $G^{\prime}$ cannot contain a good $(d-1)$-tuple.
We also need to verify that $\varepsilon^{\prime}$ is large enough so that we are able to apply the induction hypothesis on $d$, i.e. we need to check that $\varepsilon^{\prime d-2}n^{\prime}\geq 100^{d-2}$ and $\varepsilon^{\prime d-1}m^{\prime}\geq 100^{d-1}$. This is a consequence of an easy computation since
$\varepsilon^{\prime d-2}n^{\prime}\geq\left(1-\frac{5/2}{d}\right)^{d-2}\varepsilon^{d-2}\frac{\varepsilon n}{4}\geq\frac{e^{-5/2}}{4}100^{d-1}\geq 100^{d-2}.$
In the above calculation, we have used that $\varepsilon^{d-1}n\geq 100^{d-1}$ and that $\big{(}1-\frac{5/2}{d}\big{)}^{d-2}\geq e^{-5/2}$, which holds since $\big{(}1-\frac{5/2}{d}\big{)}^{d-2}$ is a decreasing function for $d\geq 3$ with the limit $e^{-5/2}$ when $d\to\infty$. A similar computation shows that $\varepsilon^{\prime d-1}m^{\prime}\geq\big{(}1-\frac{5/2}{d}\big{)}^{d}\varepsilon^{d-1}m\geq 100^{d}\varepsilon^{-1}\big{(}1-\frac{5/2}{4}\big{)}^{4}\geq 100^{d-1}$, where we have used that $\big{(}1-\frac{5/2}{d}\big{)}^{d}$ is an increasing function of $d$, when $d\geq 3$. Having verified all the assumptions, we apply the induction hypothesis to the graph $G^{\prime}$ and conclude that
$\operatorname{rs}(G^{\prime})\geq\left(1-\frac{5/2}{d}\right)^{d-2}\left(\frac{\varepsilon}{100}\right)^{d-2}|V_{1}||B|\geq e^{-5/2}\frac{\varepsilon^{d-1}}{4\cdot 100^{d-2}}mn\geq\frac{\varepsilon^{d-1}}{100^{d-1}}mn.$
Hence, we have $\operatorname{rs}(G)\geq\operatorname{rs}(G^{\prime})\geq(\frac{\varepsilon}{100})^{d-1}mn$, which completes the proof. ∎
@@
Proving the lower bounds on $\operatorname{rs}_{d}(m,n,I)$ and $\operatorname{rs}_{d}(n,m,I)$ is completely symmetric, so we just discuss $\operatorname{rs}_{d}(m,n,I)$. If $\varepsilon>100\max\{m^{-\frac{1}{d}},n^{-\frac{1}{d-1}}\}$, let us consider any sets of $n$ points and $m$ hyperplanes, say $\mathcal{P}$ and $\mathcal{H}$. The incidence graph $G=G(\mathcal{P},\mathcal{H})$ avoids the pattern $\Pi_{d}$ by Lemma 1.4 and so $G$ contains no good $d$-tuples. Since $G$ has at least $\varepsilon mn$ edges, Lemma 4.3 implies $\operatorname{rs}(G)\geq\left(\frac{\varepsilon}{100}\right)^{d-1}mn$. ∎
5 Constructions and examples
The main goal of this section is to provide examples which show that the bounds given in Theorem 1.2 and Theorem 1.3 are tight. The main building block of all constructions in this section are subspace evasive sets. Throughout this section, we consider only finite fields $\mathbb{F}$.
@@ 5 Constructions and examples
For integers $1\leq k\leq d$, there exists $s\leq d^{d}$ such that every finite space $\mathbb{F}^{d}$ contains a $(k,s)$-subspace evasive set of size at least $|\mathbb{F}|^{d-k}$.
@@
5.1 Constructions without $K_{s,s}$
In this section, we prove the following equivalent formulation of the upper bounds in Theorem 1.2.
@@ 5.1 Constructions without $K_{s,s}$
For every dimension $d$, there exist $s=s(d)$ and $c=c(d)>0$ such that the following holds. For every $\alpha>0$ and sufficiently large integer $n$, there exists a field $\mathbb{F}$, a set of $n$ points $\mathcal{P}$ and a set of $m=\lfloor n^{\alpha}\rfloor$ hyperplanes $\mathcal{H}$ in $\mathbb{F}^{d}$ such that the incidence graph of $(\mathcal{P},\mathcal{H})$ is $K_{s,s}$-free and
$I(\mathcal{P},\mathcal{H})\geq\begin{cases}n&\mbox{ if }\alpha\in[0,\beta_{1}],\\ cm^{1-\frac{1}{t}}n&\mbox{ if }\alpha\in[\alpha_{t},\beta_{t}]\mbox{ for some }t\in\{2,\dots,d\},\\ cmn^{1-\frac{1}{d+2-t}}&\mbox{ if }\alpha\in[\beta_{t-1},\alpha_{t}]\mbox{ forsome }t\in\{2,\dots,d\},\\ m&\mbox{ if }\alpha\in[\beta_{d},\infty).\end{cases}$
@@
Observe that Lemma 5.4 implies the existence of $\mathcal{P},\mathcal{H}$ with $|\mathcal{P}|=n_{0}=\Omega_{d}(p^{t}),|\mathcal{H}|=m_{0}=\Omega_{d}(p^{d+2-t})$ and $I(P,\mathcal{H})=\Omega_{d}((m_{0}n_{0})^{1-\frac{1}{d+2}})$. This only covers a sparse set of potential values of $m$ and $n$. However, by using an approximation and a subsampling trick, one can cover all possible values of $m$ and $n$.
@@
Let us first discuss the boundary cases, i.e. when $\alpha\in[0,\beta_{1}]$ or $\alpha\in[\beta_{d},\infty)$. In this case, we show that there exist sets $\mathcal{P}$ and $\mathcal{H}$ with $|\mathcal{P}|=n,|\mathcal{H}|=m=n^{\alpha}$, $I(\mathcal{P},\mathcal{H})\geq\max\{m,n\}$ and such that $G(\mathcal{P},\mathcal{H})$ does not contain $K_{2,2}$. In the case $\alpha\in[0,\beta_{1}]$, it suffices to choose $p\geq\max\{m,n\}$ and take $\mathcal{H}$ to be $m$ parallel hyperplanes in $\mathbb{F}_{p}^{d}$ and $\mathcal{P}$ to be $n$ points on one of the hyperplanes. On the other hand, when $\alpha\in[\beta_{d},\infty)$, let $\mathcal{P}$ be $n$ collinear points and $\mathcal{H}$ be a set of $m$ hyperplanes containing exactly one of these points.
@@
First, consider the case when $\alpha_{t}\leq\alpha\leq\beta_{t}$ for some $t\in\{2,\dots,d\}$. Let $p$ be the smallest prime larger than $2(sm)^{1/t}$. By Bertrand’s postulate, we have $p\leq 4(sm)^{1/t}$. By applying Lemma 5.4 with $d+2-t$ instead of $t$, we obtain a set of points $\mathcal{P}_{0}$ and a set of hyperplanes $\mathcal{H}_{0}$ in $\mathbb{F}_{p}^{d}$ with $|\mathcal{P}_{0}|=n_{0}\geq p^{d+2-t},|\mathcal{H}_{0}|=m_{0}\geq p^{t}/s$ and $I(\mathcal{P}_{0},\mathcal{H}_{0})=\frac{1}{p}|\mathcal{P}_{0}||\mathcal{H}_{0}|$. Moreover, $G(\mathcal{P}_{0},\mathcal{H}_{0})$ does not contain $K_{s,s}$ as a subgraph.
@@
Let $\mathcal{P}$ and $\mathcal{H}$ be random subsets of $P_{0},\mathcal{H}_{0}$ containing $n$ and $m$ elements, respectively. Observe that $m_{0}\geq\frac{1}{s}p^{t}\geq m$ and $n_{0}\geq p^{d+2-t}\geq m^{\frac{d+2-t}{t}}\geq m^{1/\alpha}=n$, so it indeed makes sense to consider such subsets of $\mathcal{P}_{0},\mathcal{H}_{0}$. The expected number of incidences between $\mathcal{P}$ and $\mathcal{H}$ is simply $\mathbb{E}[I(P,\mathcal{H})]=\frac{m}{m_{0}}\cdot\frac{n}{n_{0}}I(P_{0},\mathcal{H}_{0})\geq\frac{1}{p}mn$. Since $p\leq 4(sm)^{-1/t}$, picking a pair $(\mathcal{P},\mathcal{H})$ with at least as many incidences as the expectation, we get
$I(\mathcal{P},\mathcal{H})\geq\frac{1}{p}mn\geq\frac{1}{4s^{1/t}}m^{-\frac{1}{t}}\cdot mn.$
@@
In the case $\beta_{t-1}\leq\alpha\leq\alpha_{t}$, we proceed similarly. We pick $p$ to be the smallest prime larger than $2s^{\frac{1}{t}}n^{\frac{1}{d+2-t}}$, then $p<4s^{\frac{1}{t}}n^{\frac{1}{d+2-t}}$. We apply Lemma 5.4 with $d+2-t$ instead of $t$ to find a set of points $\mathcal{P}_{0}$ and a set of hyperplanes $\mathcal{H}_{0}$ in $\mathbb{F}_{p}^{d}$ with $|\mathcal{P}_{0}|=n_{0}\geq p^{d+2-t},|\mathcal{H}_{0}|=m_{0}\geq p^{t}/s$ and $I(\mathcal{P}_{0},\mathcal{H}_{0})=\frac{1}{p}|\mathcal{P}_{0}||\mathcal{H}_{0}|$. By an almost identical computation as before one can check that $n_{0}\geq n$ and $m_{0}\geq n^{\alpha}\geq m$. Hence, if $\mathcal{P}\subset\mathcal{P}_{0}$ is a random $n$-element subset and $\mathcal{H}\subset\mathcal{H}_{0}$ is a random $m$-element subset, they satisfy $\mathbb{E}(I(\mathcal{P},\mathcal{H}))\geq\frac{1}{p}mn\geq\frac{1}{4s^{1/t}}n^{-\frac{1}{d+2-t}}\cdot mn$. Therefore, we can pick a pair $(\mathcal{P},\mathcal{H})$ with at least as many incidences as the expectation, completing the proof. ∎
@@
5.2 Constructions with no large complete bipartite subgraphs

### v2

@@ 3.2 Large complete bipartite subgraph of dense incidence graphs
The last inequality is true due to simple calculations. We should also verify that $G^{\prime}$ satisfies the conditions of Lemma 3.6 by showing that $\varepsilon^{\prime d-1}m^{\prime}\geq 100^{d-1}$ and $\varepsilon^{\prime d}n^{\prime}\geq 100^{d}$. But this also follows, since we have already shown that $\varepsilon^{\prime d-1}m^{\prime}n^{\prime}\geq\varepsilon^{d-1}mn$, which implies both $\varepsilon^{\prime d-1}m^{\prime}\geq\varepsilon^{d-1}m\geq 100^{d-1}$ and $\varepsilon^{\prime d}n^{\prime}\geq\varepsilon^{d}n\geq 100^{d}$ (since $\varepsilon^{\prime d-1}m^{\prime}n^{\prime}\geq\varepsilon^{d-1}mn$ also implies $\varepsilon^{\prime}\geq\varepsilon$). We conclude that $\operatorname{rs}(G)\geq(\varepsilon/100)^{d-1}mn$, a contradiction.
[…]
Figure 2: An illustration for the proof of Lemma 3.6
@@
In what follows, we assume that $G$ contains no vertex $v\in A$ of degree less than $\left(1-\frac{2}{d}\right)\varepsilon n$ and no vertex $v\in B$ of degree less than $\left(1-\frac{2}{d}\right)\varepsilon m$. We choose a good $(d-1)$-tuple $(v_{1},\dots,v_{d-1})$ with the maximal size of the common neighbourhood, which we denote by $X=N(v_{1},\dots,v_{d-1})$. By Lemma 3.8, there must be a good $(d-1)$-tuple with $|X|\geq\left(\frac{\varepsilon}{6}\right)^{d-1}|B|$. If $A$ does not contain a good $d$-tuple in $G$, then any vertex $v\in A$ either has $X\subseteq N(v)$ or $|N(v)\cap X|=1$. Lemma 3.7 applied with $\gamma=(\frac{\varepsilon}{6})^{d-2}$ shows that $|X|\leq\left(\frac{\varepsilon}{6}\right)^{d-2}|B|$. We conclude that one can partition $A$ into two sets, $V_{1}$ and $V_{2}$, where $V_{1}=\{v\in A:X\subseteq N(v)\}$ and $V_{2}=\{v\in A:|X\cap N(v)|\leq 1\}$.
@@
Let us consider the edges incident to $X$. By our assumption on the minimum degree of vertices of $X$, we have $e(X,A)\geq|X|\cdot\left(1-\frac{2}{d}\right)\varepsilon m\geq\frac{\varepsilon}{3}m|X|$. On the other hand,
$\frac{\varepsilon}{3}m|X|\leq e(X,A)=e(X,V_{1})+e(X,V_{2})\leq|X||V_{1}|+|V_{2}|\leq|X||V_{1}|+m.$
@@
Since $|X|\geq\left(\frac{\varepsilon}{6}\right)^{d-1}n$ and $(\frac{\varepsilon}{100})^{d}n\geq 1$, we have $|X|\geq\frac{12}{\varepsilon}$. Thus, $\frac{\varepsilon}{12}m|X|\geq m$ and so $\frac{\varepsilon}{4}m|X|\leq|X||V_{1}|$, showing that $|V_{1}|\geq\frac{\varepsilon}{4}m$.
@@
since $6^{d-2}\geq 2d$ for all $d\geq 3$. Thus, the sizes of the vertex classes of $G^{\prime}$ are $m^{\prime}=|V_{1}|\geq\frac{\varepsilon m}{4}$ and $n^{\prime}=|B\backslash X|\geq n-(\frac{\varepsilon}{6})^{d-2}n\geq\frac{5}{6}n$. Finally, the above computation shows that $e(G^{\prime})=\varepsilon^{\prime}m^{\prime}n^{\prime}$, where $\varepsilon^{\prime}\geq(1-\frac{5/2}{d})\varepsilon$.
Next, we argue that $G^{\prime}$ does not have a good $(d-1)$-tuple. Suppose for contradiction that $G^{\prime}$ contains a good $(d-1)$-tuple $(u_{1},\dots,u_{d-1})$. The vertices $(u_{1},\dots,u_{d-1})$ form a good $(d-1)$-tuple in the graph $G$ as well, and their common neighbourhood contains both $X$ and at least two vertices from $B\backslash X$ (since this is a good $(d-1)$-tuple in $G^{\prime}$). But this implies $u_{1},\dots,u_{d-1}$ is a $(d-1)$-tuple with a larger common neighbourhood in $G$ than $v_{1},\dots v_{d-1}$, contradicting the maximality. Thus, we conclude that $G^{\prime}$ cannot contain a good $(d-1)$-tuple.
Furthermore, we show that $\varepsilon^{\prime d-2}m^{\prime}\geq 100^{d-2}$ and $\varepsilon^{\prime d-1}n^{\prime}\geq 100^{d-1}$. This is a consequence of an easy computation since
$\varepsilon^{\prime d-2}m^{\prime}\geq\left(1-\frac{5/2}{d}\right)^{d-2}\varepsilon^{d-2}\frac{\varepsilon m}{4}\geq\frac{e^{-5/2}}{4}100^{d-1}\geq 100^{d-2}.$
In the above calculation, we used that $\varepsilon^{d-1}n\geq 100^{d-1}$ and that $\big{(}1-\frac{5/2}{d}\big{)}^{d-2}\geq e^{-5/2}$, which holds since $\big{(}1-\frac{5/2}{d}\big{)}^{d-2}$ is a decreasing function for $d\geq 3$ with the limit $e^{-5/2}$ when $d\to\infty$. A similar computation shows that $\varepsilon^{\prime d-1}n^{\prime}\geq\big{(}1-\frac{5/2}{d}\big{)}^{d-1}\varepsilon^{d-1}\cdot\frac{5}{6}n\geq 100^{d}\varepsilon^{-1}\big{(}1-\frac{5/2}{3}\big{)}^{2}\cdot\frac{5}{6}\geq 100^{d-1}$, where we have used that $\big{(}1-\frac{5/2}{d}\big{)}^{d-1}$ is an increasing function of $d$, when $d\geq 3$. As $G$ is a counterexample with $d$ minimal, $G^{\prime}$ is not a counterexample, so
$\operatorname{rs}(G^{\prime})\geq\left(1-\frac{5/2}{d}\right)^{d-2}\left(\frac{\varepsilon}{100}\right)^{d-2}|V_{1}||B\backslash X|\geq e^{-5/2}\frac{\varepsilon^{d-1}}{4\cdot 100^{d-2}}m\cdot\frac{5}{6}n\geq\frac{\varepsilon^{d-1}}{100^{d-1}}mn.$
Hence, we have $\operatorname{rs}(G)\geq\operatorname{rs}(G^{\prime})\geq(\frac{\varepsilon}{100})^{d-1}mn$, which is a contradiction. ∎
@@
Proving the lower bounds on $\operatorname{rs}_{d}(m,n,I)$ and $\operatorname{rs}_{d}(n,m,I)$ is completely symmetric, so we just discuss $\operatorname{rs}_{d}(m,n,I)$. If $\varepsilon>100\max\{m^{-\frac{1}{d-1}},n^{-\frac{1}{d}}\}$, let us consider any sets of $m$ points and $n$ hyperplanes, say $\mathcal{P}$ and $\mathcal{H}$. The incidence graph $G=G(\mathcal{P},\mathcal{H})$ avoids the pattern $\Pi_{d}$ by Lemma 2.1 and so $G$ contains no good $d$-tuples. Since $G$ has at least $\varepsilon mn$ edges, Lemma 3.6 implies $\operatorname{rs}(G)\geq\left(\frac{\varepsilon}{100}\right)^{d-1}mn$. ∎
3.3 Constructions $K_{s,s}$-free incidence graphs
The main goal of this section is to provide examples which show that the bounds given in Theorem 1.2 are tight. The main building block of all constructions in this section are subspace evasive sets. Throughout this section, we consider only finite fields $\mathbb{F}$.
@@ 3.3 Constructions $K_{s,s}$-free incidence graphs
For integers $1\leq k\leq d$, there exists $s\leq d^{d}$ such that for every large enough finite field $\mathbb{F}$, the space $\mathbb{F}^{d}$ contains a $(k,s)$-subspace evasive set of size at least $|\mathbb{F}|^{d-k}$.
@@
We prove the following equivalent formulation of the upper bounds in Theorem 1.2.
@@
For every dimension $d$, there exist $s=s(d)$ and $c=c(d)>0$ such that the following holds. For every $\alpha>0$ and sufficiently large integer $m$, there exists a field $\mathbb{F}$, a set of $m$ points $\mathcal{P}$ and a set of $n=\lfloor m^{\alpha}\rfloor$ hyperplanes $\mathcal{H}$ in $\mathbb{F}^{d}$ such that the incidence graph of $(\mathcal{P},\mathcal{H})$ is $K_{s,s}$-free and
$I(\mathcal{P},\mathcal{H})\geq\begin{cases}m&\mbox{ if }\alpha\in(0,\beta_{1}],\\ cmn^{1-\frac{1}{t}}&\mbox{ if }\alpha\in[\alpha_{t},\beta_{t}]\mbox{ for some }t\in\{2,\dots,d\},\\ cm^{1-\frac{1}{d+2-t}}n&\mbox{ if }\alpha\in[\beta_{t-1},\alpha_{t}]\mbox{ forsome }t\in\{2,\dots,d\},\\ n&\mbox{ if }\alpha\in[\beta_{d},\infty).\end{cases}$
@@
Observe that Lemma 3.12 implies the existence of $\mathcal{P},\mathcal{H}$ with $|\mathcal{P}|=m_{0}=\Omega_{d}(p^{t}),|\mathcal{H}|=n_{0}=\Omega_{d}(p^{d+2-t})$ and $I(\mathcal{P},\mathcal{H})=\Omega_{d}((m_{0}n_{0})^{1-\frac{1}{d+2}})$. This only covers a sparse set of potential values of $m$ and $n$. However, by using an approximation and a subsampling trick, one can cover all possible values of $m$ and $n$.
@@
Let us first discuss the boundary cases, i.e. when $\alpha\in(0,\beta_{1}]$ or $\alpha\in[\beta_{d},\infty)$. In this case, we show that there exist sets $\mathcal{P}$ and $\mathcal{H}$ with $|\mathcal{P}|=m,|\mathcal{H}|=n=m^{\alpha}$, $I(\mathcal{P},\mathcal{H})\geq\max\{m,n\}$ and such that $G(\mathcal{P},\mathcal{H})$ does not contain $K_{2,2}$. In the case $\alpha\in[0,\beta_{1}]$, it suffices to choose $p\geq\max\{m,n\}$ and take $\mathcal{H}$ to be $n$ parallel hyperplanes in $\mathbb{F}_{p}^{d}$ and $\mathcal{P}$ to be $m$ points on one of the hyperplanes. On the other hand, when $\alpha\in[\beta_{d},\infty)$, let $\mathcal{P}$ be $m$ collinear points and $\mathcal{H}$ be a set of $n$ hyperplanes containing exactly one of these points.
@@
First, consider the case when $\alpha_{t}\leq\alpha\leq\beta_{t}$ for some $t\in\{2,\dots,d\}$. Let $p$ be the smallest prime larger than $2(sn)^{1/t}$. By Bertrand’s postulate, we have $p\leq 4(sn)^{1/t}$. By applying Lemma 3.12 with $d+2-t$ instead of $t$, we obtain a set of points $\mathcal{P}_{0}$ and a set of hyperplanes $\mathcal{H}_{0}$ in $\mathbb{F}_{p}^{d}$ with $|\mathcal{P}_{0}|=m_{0}\geq p^{d+2-t},|\mathcal{H}_{0}|=n_{0}\geq p^{t}/s$ and $I(\mathcal{P}_{0},\mathcal{H}_{0})=\frac{1}{p}|\mathcal{P}_{0}||\mathcal{H}_{0}|$. Moreover, $G(\mathcal{P}_{0},\mathcal{H}_{0})$ does not contain $K_{s,s}$ as a subgraph.
@@
Let $\mathcal{P}$ and $\mathcal{H}$ be random subsets of $\mathcal{P}_{0},\mathcal{H}_{0}$ containing $m$ and $n$ elements, respectively. Observe that $n_{0}\geq\frac{1}{s}p^{t}\geq n$ and $m_{0}\geq p^{d+2-t}\geq m^{\frac{d+2-t}{t}}\geq n^{1/\alpha}=m$, so it indeed makes sense to consider such subsets of $\mathcal{P}_{0},\mathcal{H}_{0}$. The expected number of incidences between $\mathcal{P}$ and $\mathcal{H}$ is simply $\mathbb{E}[I(\mathcal{P},\mathcal{H})]=\frac{m}{m_{0}}\cdot\frac{n}{n_{0}}I(\mathcal{P}_{0},\mathcal{H}_{0})\geq\frac{1}{p}mn$. Since $p\leq 4(sn)^{-1/t}$, picking a pair $(\mathcal{P},\mathcal{H})$ with at least as many incidences as the expectation, we get
$I(\mathcal{P},\mathcal{H})\geq\frac{1}{p}mn\geq\frac{1}{4s^{1/t}}n^{-\frac{1}{t}}\cdot mn.$
@@
In the case $\beta_{t-1}\leq\alpha\leq\alpha_{t}$, we proceed similarly. We pick $p$ to be the smallest prime larger than $2s^{\frac{1}{t}}m^{\frac{1}{d+2-t}}$, then $p<4s^{\frac{1}{t}}m^{\frac{1}{d+2-t}}$. We apply Lemma 3.12 with $d+2-t$ instead of $t$ to find a set of points $\mathcal{P}_{0}$ and a set of hyperplanes $\mathcal{H}_{0}$ in $\mathbb{F}_{p}^{d}$ with $|\mathcal{P}_{0}|=m_{0}\geq p^{d+2-t},|\mathcal{H}_{0}|=n_{0}\geq p^{t}/s$ and $I(\mathcal{P}_{0},\mathcal{H}_{0})=\frac{1}{p}|\mathcal{P}_{0}||\mathcal{H}_{0}|$. By an almost identical computation as before one can check that $m_{0}\geq m$ and $n_{0}\geq m^{\alpha}\geq n$. Hence, if $\mathcal{P}\subset\mathcal{P}_{0}$ is a random $m$-element subset and $\mathcal{H}\subset\mathcal{H}_{0}$ is a random $n$-element subset, they satisfy $\mathbb{E}(I(\mathcal{P},\mathcal{H}))\geq\frac{1}{p}mn\geq\frac{1}{4s^{1/t}}m^{-\frac{1}{d+2-t}}\cdot mn$. Therefore, we can pick a pair $(\mathcal{P},\mathcal{H})$ with at least as many incidences as the expectation, completing the proof. ∎
@@
3.4 Constructions of dense incidence graphs with no large bipartite subgraphs